# Isolated Causal Effects of Natural Language

### v1

@@ Abstract
As language technologies become widespread, it is important to understand how variations in language affect reader perceptions—formalized as the isolated causal effect of some focal language-encoded intervention on an external outcome. A core challenge of estimating isolated effects is the need to approximate all non-focal language outside of the intervention. In this paper, we introduce a formal estimation framework for isolated causal effects and explore how different approximations of non-focal language impact effect estimates. Drawing on the principle of omitted variable bias, we present metrics for evaluating the quality of isolated effect estimation and non-focal language approximation along the axes of fidelity and overlap. In experiments on semi-synthetic and real-world data, we validate the ability of our framework to recover ground truth isolated effects, and we demonstrate the utility of our proposed metrics as measures of quality for both isolated effect estimates and non-focal language approximations.
@@ 1 INTRODUCTION
The widespread use of language technologies has given rise to an ever-expanding amount of human- and machine-generated text data. From this vast body of data emerges the opportunity to understand how information contained in language relates to real-world outcomes. For example, can we learn what makes a helpful product review? Is it the positive emotion in the review, or how engaging the review is to read?
(a) The isolated causal effect of a language attribute like netspeak can be learned by decomposing a text $X$ into focal language $a(X)$ and non-focal language $a^{\mathsf{c}}(X)$.
(b) Non-focal language cannot be measured directly and instead must be approximated. Which approximation we choose can significantly affect the isolated effect estimate.
@@
Figure 1: Isolated causal effects allow us to understand how choices of language affect reader perceptions.
The ways that these language choices affect reader perceptions can be formalized as the isolated causal effect of some focal language-encoded intervention—often a linguistic attribute—on an external outcome (Figure 1(a)). Because the effect of such an intervention may depend on the surrounding linguistic context, we define an isolated causal effect (or isolated effect for brevity) as the average causal effect of the focal text intervention over all possible variations of the rest of the text.
Estimating an isolated causal effect poses two major challenges. (1) We must be able to define in formal terms and approximate all variance contained in the non-focal language: that is, all parts of the language external to the intervention of interest. (2) The way we choose to model the non-focal language affects our isolated effect estimate, so we must be able to evaluate both the quality of its approximation and the quality of the isolated effect estimate itself.
@@
In this paper, we introduce a formal estimation framework for isolated causal effects. Within this framework, we explore how the way we approximate non-focal language impacts isolated effect estimates (Figure 1(b)), and we present two metrics for evaluating the quality of non-focal language approximations and isolated effect estimates along the axes of fidelity and overlap, along with a third summary metric. To evaluate along these arms, we draw on the statistical principle of omitted variable bias (OVB), a measure of bias caused by the omission of relevant variables.
@@
In our experiments, [^1] Our code is publicly available at https://github.com/torylin/isolated-text-effects. we demonstrate the validity of our effect estimation method and metrics on both semi-synthetic and real-world data. Using settings where we know the ground truth, we observe that our estimation framework is able to recover the true isolated effect across a variety of interventions. We further show that our metrics correspond closely to how well an estimator is able to recover the true effect, while also providing additional insight into why an estimate is or is not correct. As a result, we propose that these metrics may be particularly useful in real-world settings where the true isolated effect is unknown.
@@ 2 PROBLEM SETTING
Consider a text dataset $D=\{(X_{1},Y_{1}),\dots,(X_{n},Y_{n})\}$ where texts $X_{i}\in\mathcal{X}$ are drawn i.i.d. from a distribution $P$, and individuals with potential outcome functions $Y_{i}(\cdot):\mathcal{X}\rightarrow\mathbb{R}$ are drawn i.i.d. from the population $\mathcal{G}$. We assume that texts are randomly assigned to individuals, as is very often the case for natural language datasets—a primary example being crowdsourced datasets (Lin et al.,, 2024). Under random assignment, all confounding is contained within the text.
Now let $X$ be parameterized as $X=\{a(X),a^{\mathsf{c}}(X)\}$, where $a(\cdot):\mathcal{X}\rightarrow\{0,1\}$ is the intervention—the focal language attribute for which we want to learn a causal effect—and $a^{\mathsf{c}}(\cdot):\mathcal{X}\rightarrow\mathbb{R}^{d}$ is the non-focal portion of the text. $a(\cdot)$ is also known as a “codebook function” (Egami et al.,, 2022).
@@
Definition 1 (Isolated causal effect).
@@
$\tau^{*}$ $=E_{Y(\cdot)\sim\mathcal{G}}[\mathbb{E}_{a^{\mathsf{c}}(X)^{*}\sim P^{*}}[Y(a(X)=1,a^{\mathsf{c}}(X)^{*})$
@@
We distinguish this from the natural causal effect defined by Lin et al., (2023), where $a^{\mathsf{c}}(X)$ follows its natural distribution for both conditions of $a(X)$. The natural causal effect is the collective effect of the focal language attribute $a(X)$ and the parts of the non-focal language with which it is naturally correlated.
We invoke three assumptions for valid causal inference:
@@
In practice, it is reasonable to believe that the overlap assumption is fulfilled since any representation of the non-focal language $a^{c}(X)$ is non-exhaustive. Under such circumstances, $a(X)$ cannot be determined solely from $a^{c}(X)$.
@@ 3 ISOLATED CAUSAL EFFECTS OF LANGUAGE
In this section, we describe how to identify, estimate, and evaluate the quality of isolated effects of language. We define estimands for the isolated average treatment effect and isolated average treatment effect on the treated; present doubly robust estimators for each; and discuss how approximating non-focal language during estimation can give rise to omitted variable bias. Derivations and technical results are in Appendix 1.
@@ 3.1 Identification
The definition of the isolated causal effect requires that $a^{\mathsf{c}}(X)$ follow the same target distibution $P^{*}$ when $a(X)=1$ and when $a(X)=0$. To induce $a^{\mathsf{c}}(X)$ to follow a specific target distribution, we draw on importance weighting (IPW) principles to transport $a^{\mathsf{c}}(X)$ from its natural distribution to the target distribution $P^{*}$, then supplement this with an outcome model.
First, let the importance weight $\gamma$ be defined as:
@@
Next, let the outcome model $g$ be defined as follows:
$g(a^{\prime},a^{\mathsf{c}}(X))=E_{Y(\cdot)\sim\mathcal{G}}[Y(a^{\prime},a^{\mathsf{c}}(X))]$
Finally, we identify the estimand $\tau^{*}$ from the observed data $D=(X,Y)$, where $X\sim P$ and $Y$ follows the resulting induced distribution on the observed responses $P_{y}$. Given access to the data-generating process of $P^{*}$, i.e., assuming we have access to texts $X^{*}\sim P^{*}$, we have that the isolated effect is
@@
This doubly robust construction—standard in many causal inference problems—confers robustness to misspecification or mis-estimation in either the IPW term or the outcome modeling term (Robins et al.,, 1994).
While $P^{*}$ can be any distribution over $a^{\mathsf{c}}(X)$ (discussed further in Appendix 1.1.3), there are two important cases of $P^{*}$ on which we focus. First, we can set $P^{*}(a^{\mathsf{c}}(X))=P(a^{\mathsf{c}}(X))$, where again $P$ is the distribution of the observed $X$. We refer to the isolated effect in this case as the Isolated Average Treatment Effect (IATE) in the corpus from which the texts originate. The corresponding importance weight becomes:
@@ 3.2 Estimation
Nuisance parameters.
To estimate $\tau_{DR}$, several nuisance parameters need to first be estimated: the outcome model $g$ and the importance weight $\gamma$. Estimating $\gamma$ further requires both estimating the probability $P(a(X)=a^{\prime}|a^{\mathsf{c}}(X))$ and approximating the non-focal language $a^{\mathsf{c}}(X)$ with a language representation.
We refer to the approximation of the non-focal language using the notation $a_{s}^{\mathsf{c}}(X)$, where the $s$ subscript indicates that this is a mapping of the high-dimensional non-focal language $a^{\mathsf{c}}(X)$ to a lower-dimensional “short” representation space $\mathbb{R}^{d}$ (following the terminology used to denote a smaller feature set in Chernozhukov et al., (2024)). With this mapping, a classifier can be trained on a separate sample to predict $a(X)$ given $a_{s}^{\mathsf{c}}(X)$ as input. Such a classifier outputs predicted probabilities $\widehat{P}(a(X)=a^{\prime}|a_{s}^{\mathsf{c}}(X))$, which can be used to estimate $\widehat{\gamma}$.
Using the approximation $a_{s}^{\mathsf{c}}(X)$, an outcome model $\widehat{g}(a(X),a_{s}^{\mathsf{c}}(X))$ can also be estimated on an separate sample where both texts and outcomes are available.
Estimator.
Consider data $D=(X_{i},Y_{i})$, $X_{i}\sim P$ and $Y_{i}\sim P_{y}$; and $X_{j}\sim P^{*}$ ($i\in[n],j\in[m]$). Then the estimator for $\tau^{*}$ is given by
@@
where the estimated $\hat{\gamma}$ uses the appropriate probability estimates from the IATE and IATT definitions above.
Like other doubly robust estimators, $\tau_{DR}$ has a number of desirable properties. First, as long as either the weights $\gamma$ or the outcome model $g$ are correct—i.e., $\widehat{\gamma}=\gamma$ or $\widehat{g}=g$—then $\tau_{DR}$ is an unbiased estimator for $\tau^{*}$. Moreover, the estimator is asymptotically normal with a closed-form variance, allowing for estimation of trustworthy confidence intervals. See Kennedy, (2023) for a review on these types of estimators.
3.3 The Fidelity-Overlap Tradeoff
Omitted variable bias.
When representing natural language, including all “variables” in modeling or estimation is not feasible, as a full representation of language is nearly infinitely high-dimensional (consider a one-hot encoding vector where each column is one word, with the total dimensionality being the size of the vocabulary of the English language). Instead, the non-focal language is more realistically approximated as the “short” (i.e., lower-dimensional) representation $a_{s}^{\mathsf{c}}(X)$. However, representations of language necessarily omit information relative to the original text.
In this section, we link the notion of information loss from language representation to the concept of omitted variable bias (OVB). We use recent work on establishing OVB bounds for non-parametric models (Chernozhukov et al.,, 2024) and adapt it to a natural language setting to study the impact of omitted information in approximations of non-focal language. We consider OVB to be responsible for a large portion of the estimation error in recovering isolated causal effects.
Fidelity and overlap.
When representing non-focal language, two primary factors contribute to OVB. The first is the fidelity of the “short” outcome model $g(a^{\prime},a_{s}^{\mathsf{c}}(X))$, or how close the short model is to the true outcome model $g(a^{\prime},a^{\mathsf{c}}(X))$. The second is the overlap of the “short” importance weight $\gamma(a^{\prime},a_{s}^{\mathsf{c}}(X))$. Overlap violations can occur if the $a(X)$ classifier can fully predict $a(X)$ from $a_{s}^{c}(X)$, as the predicted probabilities $\widehat{P}(a(X)=a^{\prime}|a_{s}^{\mathsf{c}}(X))$ will either be 0 or 1 (this could occur, for instance, if $a(X)$ were contained in $a_{s}^{c}(X)$). While we assume in Section 2 that overlap is fulfilled for the true $P(a(X)=a^{\prime}|a^{\mathsf{c}}(X))$, approximation and estimation may still introduce overlap problems. Moreover, probabilities very close to 0 and 1 (“near overlap violations”) can still skew the importance weights $\gamma$ to extreme values. Consequently, even slight differences between the approximated and true non-focal language can have large impacts on the estimated isolated effect $\widehat{\tau}_{DR}$.
A fidelity-overlap tradeoff emerges when choosing how to approximate the non-focal language $a^{\mathsf{c}}(X)$ as $a_{s}^{\mathsf{c}}(X)$. If $a_{s}^{\mathsf{c}}(X)$ is a high-dimensional dense representation like a language model embedding, then model fidelity is likely to be good, as the short outcome model $g(a(X),a_{s}^{c}(X))$ has plenty of information with which to make predictions. However, $P(a(X)=a^{\prime}|a_{s}^{c}(X))$ is also more likely to be close to 0 or 1, as the greater number of dimensions makes it more likely that certain values of $a_{s}^{\mathsf{c}}(X)$ are almost exclusively seen with either $a(X)=1$ or $a(X)=0$.
We note that compared to the IATE, the IATT is less susceptible to overlap violations, as the IATT requires only that $P(a(X)=1|a_{s}^{\mathsf{c}}(X))<1$.
Metrics.
To choose a non-focal language approximation that strikes the optimal balance between model fidelity and overlap, we present metrics derived from the operationalization of OVB described by Chernozhukov et al., (2024). The fidelity metric $\sigma^{2}$ and the overlap metric $\nu^{2}$ are identified as:

### v3

@@ Abstract
As language technologies become widespread, it is important to understand how changes in language affect reader perceptions and behaviors. These relationships may be formalized as the isolated causal effect of some focal language-encoded intervention (e.g., factual inaccuracies) on an external outcome (e.g., readers’ beliefs). In this paper, we introduce a formal estimation framework for isolated causal effects of language. We show that a core challenge of estimating isolated effects is the need to approximate all non-focal language outside of the intervention. Drawing on the principle of omitted variable bias, we provide measures for evaluating the quality of both non-focal language approximations and isolated effect estimates themselves. We find that poor approximation of non-focal language can lead to bias in the corresponding isolated effect estimates due to omission of relevant variables, and we show how to assess the sensitivity of effect estimates to such bias along the two key axes of fidelity and overlap. In experiments on semi-synthetic and real-world data, we validate the ability of our framework to correctly recover isolated effects and demonstrate the utility of our proposed measures.
causal inference, natural language processing, omitted variable bias
@@ 1 Introduction
The widespread use of language technologies has given rise to an ever-expanding amount of human- and machine-generated text data. From this vast body of data emerges the opportunity to understand how information contained in language relates to real-world outcomes. Elucidating these relationships can help answer scientifically interesting questions and provide interpretability to texts and the models that generate them. For instance, what language attributes (e.g., profanity) cause readers to perceive a passage of text as hateful? Does use of rapport-building language by therapists help to improve patients’ mental health outcomes? Do factual inaccuracies propagated in machine-generated texts have negative impacts on readers’ beliefs or behaviors?
@@
Figure 1: Isolated causal effects allow us to understand how changes in language affect reader perceptions and behaviors.
The way in which language choices affect reader perceptions can be formalized as the causal effect of some language-encoded intervention—often a linguistic attribute—on an external outcome (Lin et al., 2023). However, because language is highly aliased (i.e., correlated with itself), the effect of such an intervention may be influenced by the surrounding linguistic context (Fong & Grimmer, 2023). For instance, machine-generated texts with factual inaccuracies may also contain other undesirable attributes likely to influence readers’ reactions, such as inflammatory language. If the causal effect of the factual inaccuracies is estimated without accounting for aliased attributes, the resulting estimate may contain the collective effect of both the factual inaccuracies and some portion of the related inflammatory language—making it difficult to determine whether action should be taken to address factual inaccuracies only, inflammatory language only, or both.
Motivated by this limitation, we propose a new target of inference: the isolated causal effect for natural language (or isolated effect for brevity). We define the isolated effect as the causal effect of only the part of the language contained in the intervention, or the average causal effect of the focal text intervention over all possible variations of the rest of the text (Figure 1(a)).
In practice, estimating such an effect poses several major challenges. (1) We must be able to formally define and approximate not only the focal intervention but also the non-focal language of a text: that is, all parts of the text external to the focal intervention. (2) Incorrect modeling of the non-focal language can lead to a biased or invalid estimate of the isolated effect due to omission of key language context—a form of omitted variable bias (OVB) (Figure 1(b)). In other words, valid isolated causal effects can only be measured if the non-focal language is well-approximated. Therefore, it is important to be able to assess the robustness of the isolated effect estimate to errors or omissions in the non-focal language approximations.
@@
To address these challenges and to provide a practical path toward estimating isolated effects, this paper introduces a formal estimation framework for isolated effects of language. Within this framework, we explore how the way we approximate non-focal language impacts isolated effect estimates due to omission of key variables, and we draw on OVB principles to define measures that assess the sensitivity of effect estimates to bias along the two key axes of fidelity and overlap.
@@
Our experiments [^1] Our data and code are publicly available at https://github.com/torylin/isolated-text-effects. demonstrate the validity of our framework on both semi-synthetic and real-world data. Using evaluation settings where the ground truth is known, we observe that our estimation framework is able to recover the true isolated effect across multiple interventions. We further show that our measures of overall robustness to OVB correspond closely to how well an estimator is able to recover the true effect, while fidelity and overlap provide additional insight into why an estimate is or is not correct. We suggest that these measures may be particularly useful for analysis in real-world settings where the true isolated effect is unknown.
@@ 2 Problem Setting
Consider a text dataset $D=\{(X_{1},Y_{1}),\dots,(X_{n},Y_{n})\}$ where texts $X_{i}\in\mathcal{X}$ are drawn i.i.d. from a distribution $P$, and individuals with potential outcome functions $Y_{i}(\cdot):\mathcal{X}\rightarrow\mathbb{R}$ are drawn i.i.d. from the population $\mathcal{G}$, where $Y_{i}(x)$ denotes the potential outcome of individual $i$ if they were to read text $x$ (Neyman, 1923 [1990]; Rubin, 1974). We study a setting in which all confounding between the intervention and the outcome is captured in the text. Such settings are an important and common case in natural language processing (NLP) due to the widespread practice of labeling text data using external annotators who are randomly assigned to texts. This assignment mechanism in effect creates a randomized text experiment, eliminating confounding external to the text (Lin et al., 2024). Prominent NLP benchmark datasets such as SST (Socher et al., 2013), SQuAD (Rajpurkar et al., 2018), and MNLI (Williams et al., 2018), for instance, all fall under this category.
Now let $X$ be parameterized as $X=\{a(X),a^{\mathsf{c}}(X)\}$, where $a(\cdot):\mathcal{X}\rightarrow\{0,1\}$ is the intervention—the focal language attribute for which we learn a causal effect—and $a^{\mathsf{c}}(\cdot):\mathcal{X}\rightarrow\mathbb{R}^{d}$ is the non-focal portion of the text. In this setting, we consider $a(\cdot)$ to be a known mapping from the text to the intervention of interest. $a(\cdot)$ is also known as a “codebook function” (Egami et al., 2022)
@@
Definition 2.1 (Isolated causal effect).
@@
$\tau^{*}$ $=\mathbb{E}_{Y(\cdot)\sim\mathcal{G}}[\mathbb{E}_{a^{\mathsf{c}}(X)^{*}\sim P^{*}}[Y(a(X)=1,a^{\mathsf{c}}(X)^{*})$
@@
We distinguish this from the natural causal effect defined by Lin et al. (2023), where $a^{\mathsf{c}}(X)$ follows its natural distribution for both conditions of $a(X)$. The natural causal effect is the collective effect of the focal language attribute $a(X)$ and the parts of the non-focal language with which it is naturally correlated.
We generalize three common assumptions for valid causal inference to the language setting:
@@
As we mention earlier, the assumption of no unmeasured confounding is commonly fulfilled for NLP datasets due to text-annotator assignment protocols that eliminate external confounding. In practice, it is also reasonable to believe that the overlap assumption is fulfilled since any representation of the non-focal language $a^{c}(X)$ is non-exhaustive, and so $a(X)$ cannot be determined solely from the representation of $a^{c}(X)$. The most difficult of the three assumptions to fulfill then is consistency, i.e., that observed outcomes correspond to potential outcomes. If the approximation of the non-focal language $a^{\mathsf{c}}(X)$ is missing important information, then consistency may not hold. Part of the technical contribution of this paper is to characterize the implications of this assumption failing to hold.
@@ 3 Isolated Causal Effects of Language
In this section, we describe how to identify, estimate, and evaluate the quality of isolated effects of language. We define estimands for isolated effects, present doubly robust estimators, and discuss how approximating non-focal language during estimation can give rise to omitted variable bias. Derivations and technical results are in Appendix A.
@@ 3.1 Identification
The definition of the isolated causal effect requires that $a^{\mathsf{c}}(X)$ follow the same target distribution $P^{*}$ when $a(X)=1$ and when $a(X)=0$, even if it does not do so naturally. To induce $a^{\mathsf{c}}(X)$ to follow the same specific target distribution under both conditions of $a(X)$, we draw on importance weighting (IPW) principles to transport $a^{\mathsf{c}}(X)$ from its natural distribution $P$ to the target distribution $P^{*}$, then supplement this with an outcome model.
First, let us define the transporting importance weight $\gamma$ as:
@@
Using the importance weight, we can identify the estimand $\tau^{*}$ from the observed data $D=(X,Y)$, where $X\sim P$ and $Y$ follows the resulting induced distribution on the observed responses $P_{y}$:
$\tau^{*}$ $=\mathbb{E}_{D}[\gamma(a(X),a^{\mathsf{c}}(X))Y]$
$=\mathbb{E}_{D}\left[\frac{a(X)P^{*}(a^{\mathsf{c}}(X))}{P(a^{\mathsf{c}}(X))P(a(X)=1|a^{\mathsf{c}}(X))}Y\right]$
$\quad-\mathbb{E}_{D}\left[\frac{(1-a(X))P^{*}(a^{\mathsf{c}}(X))}{P(a^{\mathsf{c}}(X))P(a(X)=0|a^{\mathsf{c}}(X))}Y\right]$
This identifies the isolated effect as a difference in importance-weighted averages of the outcome between texts with and without the focal language attribute $a(X)$.
If we use only the importance weights, however, we run the risk that errors or misspecifications in the weights will lead to errors in the estimated isolated effect. Therefore, building on causal inference principles in non-language settings, we can also incorporate an outcome model $g$, defined as:
$g(a^{\prime},a^{\mathsf{c}}(X))=\mathbb{E}_{Y(\cdot)\sim\mathcal{G}}[Y(a^{\prime},a^{\mathsf{c}}(X))].$
Assuming we have access to texts $X^{*}\sim P^{*}$ (or have access to the data-generating process of $P^{*}$), we can identify $\tau^{\ast}$ using the following doubly robust construction:
@@
This construction—commonly used in causal inference to identify and estimate unbiased effects and increasingly used in machine learning contexts as well—confers robustness to misspecification or mis-estimation in either the IPW term or the outcome modeling term (Robins et al., 1994; Byrd & Lipton, 2019; Kallus et al., 2022).
While $P^{*}$ can be any distribution over $a^{\mathsf{c}}(X)$ (discussed further in Appendix A.1.3), in practice it can be unclear how to define and estimate $P^{*}(a^{\mathsf{c}}(X))$ explicitly, as this requires characterizing a full distribution over the non-focal text probabilities. Therefore, we introduce two important realistic choices of $P^{*}$ that make the problem tractable.
First, we can set $P^{*}(a^{\mathsf{c}}(X))=P(a^{\mathsf{c}}(X))$, where again $P$ is the distribution of the observed $X$. We refer to the isolated effect in this case as the Isolated Average Treatment Effect (IATE) in the corpus from which the texts originate. The corresponding importance weight becomes:
@@ 3.2 Estimation
Nuisance parameters. To estimate $\tau_{DR}$, several nuisance parameters need to first be estimated: the outcome model $g$ and the importance weight $\gamma$. This requires approximating the non-focal language $a^{\mathsf{c}}(X)$ with a language representation. We refer to the approximation of the non-focal language using the notation $a_{s}^{\mathsf{c}}(X)$, where the $s$ subscript indicates that this is a mapping of the high-dimensional non-focal language $a^{\mathsf{c}}(X)$ to a lower-dimensional “short” representation space $\mathbb{R}^{d}$ (following the terminology used to denote a smaller feature set in Chernozhukov et al. (2024)).
With this mapping, a classifier can be trained on a separate sample to predict $a(X)$ given $a_{s}^{\mathsf{c}}(X)$ as input. Such a classifier outputs predicted probabilities $\widehat{P}(a(X)=a^{\prime}|a_{s}^{\mathsf{c}}(X))$, which can be used to estimate $\widehat{\gamma}$. Using the approximation $a_{s}^{\mathsf{c}}(X)$, an outcome model $\widehat{g}(a(X),a_{s}^{\mathsf{c}}(X))$ can also be estimated on an separate sample where both texts and outcomes are available.
Estimator. Consider data $D=(X_{i},Y_{i})$, $X_{i}\sim P$ and $Y_{i}\sim P_{y}$; and $X_{j}\sim P^{*}$ ($i\in[n],j\in[m]$). Then the estimator for $\tau^{*}$ is given by
@@
where the estimated $\widehat{\gamma}$ uses the appropriate probability estimates from the IATE and IATT definitions above.
Like other doubly robust estimators, $\tau_{DR}$ has a number of desirable properties. First, as long as either the weights $\gamma$ or the outcome model $g$ are correct—i.e., $\widehat{\gamma}=\gamma$ or $\widehat{g}=g$—then $\tau_{DR}$ is an unbiased estimator for $\tau^{*}$. Moreover, the estimator is asymptotically normal with a closed-form variance, allowing for estimation of trustworthy confidence intervals. See Kennedy (2024) for a review on these types of estimators.
3.3 Sensitivity to Omitted Variable Bias
Omitted variable bias. When representing natural language, including all “variables” in modeling is not feasible, as a full representation of language is nearly infinitely high-dimensional (e.g., a one-hot encoding of the entire English vocabulary). Instead, the non-focal language is more realistically approximated as the “short,” lower-dimensional representation $a_{s}^{\mathsf{c}}(X)$ (e.g., a language model embedding). However, representations of language necessarily omit information relative to the original text. In this section, we link the notion of information loss from language representation to omitted variable bias. We use recent work on establishing OVB bounds for non-parametric models (Chernozhukov et al., 2024) and adapt it to a natural language setting to study the impact of omitted information in approximations of non-focal language and isolated effect estimates.
We begin by defining the fidelity metric $\sigma^{2}$ and the overlap metric $\nu^{2}$: